An ideal gas absorbs 5.00 × 10^3 J energy while doing 2.00 × 10^3 J of work on the environment during a constant pressure process (a) What is the change in the internal energy of the gas?

(b) If the internal energy now drops by 4.50 × 10^3 J and 2.00 × 10^3 J is expelled from the system, find the change in volume, assuming a constant pressure process at 1.01 × 10^5 Pa.

1
Expert's answer
2022-06-06T08:54:39-0400

a)

"\u2206U=Q-A=3~kJ,"

b)

"∆V=\frac{∆U-Q}p=24.8~l."